OpenStudy (anonymous):

remote controlled car is moving in a vacant parking lot. The velocity of the car as a function of time is given by v= [5.00 m/s – (0.0180 m/s2)t^2 ]i+[2.00 m/s + (0.550 m/s2)t ]j . What are the magnitude and direction of the velocity of the car at t= 8.00 s?

OpenStudy (lgbasallote):

just substitute 8 into all the t

OpenStudy (lgbasallote):

then use the formula \[\huge magnitude = \sqrt{\hat i^2 + \hat j^2}\] \[\huge direction = \tan^{-1} \left(\frac{\hat j}{\hat i} \right)\] by i and j i mean the expressions for x and y respectively

OpenStudy (lgbasallote):

does that help?